# HoPE: Hybrid of Position Embedding for Length Generalization in Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) have made significant progress in multimodal tasks. However, their performance often deteriorates in long-context scenarios, particularly long videos. While Rotary Position Embedding (RoPE) has been widely adopted for length generalization in Large Language Models (LLMs), extending vanilla RoPE to capture the intricate spatial-temporal dependencies in videos remains an unsolved challenge. Existing methods typically allocate different frequencies within RoPE to encode 3D positional information. However, these allocation strategies mainly rely on heuristics, lacking in-depth theoretical analysis. In this paper, we first study how different allocation strategies impact the long-context capabilities of VLMs. Our analysis reveals that current multimodal RoPEs fail to reliably capture semantic similarities over extended contexts. To address this issue, we propose HoPE, a Hybrid of Position Embedding designed to improve the long-context capabilities of VLMs. HoPE introduces a hybrid frequency allocation strategy for reliable semantic modeling over arbitrarily long context, and a dynamic temporal scaling mechanism to facilitate robust learning and flexible inference across diverse context lengths. Extensive experiments across four video benchmarks on long video understanding and retrieval tasks demonstrate that HoPE consistently outperforms existing methods, confirming its effectiveness. Code is available at https://github.com/hrlics/HoPE.
@@ 1 Introduction
Vision-Language Models (VLMs) [1, 2, 3, 4, 5] have achieved remarkable success in multimodal tasks, including visual question answering [6, 7], image captioning [8, 9], and cross-modal retrieval [10, 11]. However, VLMs suffer from significant performance degradation in long-context scenarios, particularly long videos [12, 13, 14, 15]. In such settings, VLMs even struggle with simple tasks like object counting and temporal localization [16, 17], revealing a critical limitation in their ability to model extended spatial-temporal dependencies. This limitation substantially hinders their real-world deployment, where input length often exceeds the context window they have been pretrained on.
Rotary Position Embedding (RoPE) [18] has been widely adopted for length generalization in text-based LLMs [19, 20, 21]. Specifically, RoPE incorporates positional information by partitioning the query and key vectors into 2-dimensional pairs and rotating each pair at a unique frequency that decreases as the dimensional index increases. Despite its advantages, directly applying 1D RoPE fails to capture the intricate spatial-temporal dependencies in videos. Several methods have been proposed to extend 1D RoPE for multimodal inputs [2, 22, 23, 23]. Among these, the most common approach is to allocate different frequencies to encode different positional components, as shown in the upper plots of Figure 1. For example, M-RoPE [2] allocates the highest frequencies for temporal modeling ($t$), and the remaining low frequencies for spatial modeling ($x,y$). In contrast, VideoRoPE [23] proposes to allocate the lowest frequencies for temporal dimensions ($t$), and further applies a fixed scaling factor to scale the temporal indices of visual tokens, as shown in the lower plots of Figure 1. Despite their improved performance, two significant challenges remain unsolved. Firstly, current methods mainly rely on heuristics rather than theoretical analysis to determine the frequency allocation strategy. Second, applying a fixed scaling factor for all videos is suboptimal in real-world scenarios, where videos proceed with different speeds and demonstrate significant variance in information densities.
@@
To address these challenges, we first conduct an in-depth theoretical analysis of what ideal properties multimodal RoPE should hold in Section 3. Our analysis reveals that: (1) the flattening operation in vanilla RoPE inherently violates spatial-temporal locality prior, which is crucial in video modeling; (2) despite diverse frequency allocation strategies, existing multimodal RoPE variants fail to reliably capture semantic similarities over extended contexts; (3) temporal scaling of visual tokens should include both compression and expansion to accommodate varying video speeds in real-world scenarios. Guided by these insights, we propose HoPE, a Hybrid of Position Embedding designed to improve the long-context capabilities of VLMs. As shown in Figure 1, HoPE first introduces a hybrid frequency allocation strategy to facilitate long-range semantic modeling. In this strategy, higher frequencies, which are more sensitive to positional differences and better at capturing local features, are allocated for spatial positions ($x,y$) in an interleaved manner. Meanwhile, the lower frequencies are directly set to zero and allocated for temporal modeling ($t$) to enable reliable semantic modeling. Moreover, HoPE develops a dynamic temporal scaling mechanism for lengthy visual tokens. This mechanism not only enhances VLMs’ robustness to various video speeds, which are common in real-world scenarios, but also offers flexible scaling during inference across diverse context lengths.
Main Contributions. (1) To our best knowledge, we are the first to provide a theoretical analysis of how different frequency allocation strategies impact the performance of multimodal RoPE. These findings can be further utilized for the design and analysis of future multimodal RoPE variants. (2) Guided by our analysis, we propose HoPE, which consists of a hybrid frequency allocation strategy for reliable semantic modeling in long-context scenarios, and a dynamic temporal scaling mechanism for robust and flexible temporal comprehension. (3) Extensive experiments on four video benchmarks demonstrate that HoPE consistently outperforms existing RoPE variants, achieving improvements of 22.23% and 8.35% on long video retrieval and long video understanding tasks, confirming its effectiveness.
@@ 2 Preliminaries
$\mathbf{R}_{n-m}=\begin{pmatrix}\cos\theta_{0}(n-m)&-\sin\theta_{0}(n-m)&\cdots&0&0\\ \sin\theta_{0}(n-m)&\cos\theta_{0}(n-m)&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\cos\theta_{(d/2-1)}(n-m)&-\sin\theta_{(d/2-1)}(n-m)\\ 0&0&\cdots&\sin\theta_{(d/2-1)}(n-m)&\cos\theta_{(d/2-1)}(n-m)\end{pmatrix}.$ (3)
It can be observed that through pairwise attention computation, the final rotation matrix naturally incorporates the relative position information $(n-m)$ between the query-key pair.
@@ 3 Analysis
In this section, we conduct a comprehensive theoretical analysis on multimodal RoPE variants, aiming to answer the following questions: (1) Is vanilla RoPE enough for long-context VLMs? (2) How do different frequency allocation strategies impact semantic modeling in long-range multimodal contexts? (3) How should we assign the temporal indices for text and visual tokens?
@@ 3.1 Vanilla RoPE Fails in Spatial-Temporal Structure
Several recent VLMs [1, 28, 29, 15, 5] still use vanilla RoPE for multimodal inputs. In their approach, each video frame is first encoded by a vision encoder (e.g., ViT [30]) and then flattened into a sequence of patch-level tokens. These visual tokens will be treated equally as text tokens for positional encoding, with each token only incorporating 1D temporal information. We show in Proposition 3.1 that this approach, while easy to implement, distorts spatial-temporal localities and fundamentally limits VLMs’ ability to model extended spatial-temporal dependencies.
@@
Conclusion 1: Directly applying vanilla RoPE for multimodal long-context inputs inherently fails to capture their rich spatial-temporal dependencies.
@@ 3.2 Current Multimodal RoPEs Are Unreliable in Long-Range Semantic Modeling
To capture the spatial-temporal structure of multimodal inputs, a recent VLM, Qwen2-VL [2] has introduced a Multimodal Rotary Position Embedding (M-RoPE). Concretely, M-RoPE partitions the $128$-dimensional rotary embedding into three distinct groups: the first $32$ dimensions for temporal information ($t$), the subsequent $48$ dimensions for horizontal spatial information ($x$), and the final $48$ dimensions for vertical spatial information ($y$), i.e., $\mathbf{R}_{t,x,y}=\operatorname{diag}(\mathbf{R}_{t},\mathbf{R}_{x},\mathbf{R}_{y})$. While this approach realizes a naive extension for RoPE, a fundamental question remains to be answered:
@@
This question arises from the fact that in RoPE, different dimensions carry unique frequencies ($\theta_{i}=b^{-2i/d}$, $i\in\{0,1,\dots,d/2-1\}$), as shown in Equation 3. Therefore, different strategies exist for frequency allocation in multimodal RoPE. As shown in Figure 1, M-RoPE allocates the highest frequencies for $t$, intermediate frequencies for $x$, and the lowest frequencies for $y$. In contrast, VideoRoPE [23] proposes to assign the lowest frequencies for temporal modeling $(t)$ and high frequencies to spatial dimensions $(x,y)$. Their empirical justification stems from attention pattern analysis, which reveals that dimensions encoded with the lowest frequencies exhibit a more pronounced attention sink phenomenon [31], which has proven to be effective in long-context modeling. However, we argue that using the lowest frequencies for temporal modeling is still unreliable in capturing semantic similarities in extended multimodal contexts. Specifically, we first introduce semantic preference, a property where attention mechanisms should prioritize semantically similar tokens regardless of their relative distance, and formally define this concept in Definition 4.
@@
$\mathbb{E}_{\mathbf{q},\mathbf{k},\delta}[\mathbf{q}\mathbf{R}_{\Delta t\Delta x\Delta y}\mathbf{k}^{\prime\top}-\mathbf{q}\mathbf{R}_{\Delta t\Delta x\Delta y}\mathbf{k}^{\top}]\geq 0,$ (4)
@@
Let $X=[x_{1},x_{2},\dots,x_{L}]$ be an input sequence, and let RoPE use any fixed set of temporal frequencies (e.g., highest or lowest). Then there exists a critical length $L_{c}$ such that for all $L\geq L_{c}$, the semantic preference property (Definition 4) is violated.
Conclusion 2: There exists different frequency allocation strategies to extend vanilla RoPE to multimodal RoPE. However, we prove that none of these strategies can reliably maintain the semantic preference property over a sufficiently long context.
3.3 How to Assign Position Index for Multimodal Inputs?
Currently, most VLMs [1, 2, 3, 4, 14, 15] adopt the same temporal stride for video frames and text tokens, as shown in Figure 1. However, this approach overlooks the inherent difference in information densities between text and visual tokens. To address this issue, VideoRoPE [23] proposes to apply a fixed scaling factor (implemented as $2$) to adjust the temporal indices of visual tokens, achieving better empirical performance. However, this rigid scaling approach lacks the flexibility needed for diverse real-world videos, which naturally vary in pace and information density. A more ideal approach would incorporate both temporal compression and expansion capabilities, allowing the model to learn multi-scale temporal relationships, thereby enabling more robust temporal modeling.
Conclusion 3: Temporal index scaling is beneficial to balancing multimodal information. However, current methods lack flexibility and bidirectionality in temporal scaling.
4 HoPE: Hybrid of Position Embedding for Vision-Language Models
To address the above challenges, we propose HoPE, a Hybrid of Position Embedding designed to improve the long-context capability of VLMs. As illustrated in Figure 1 and Figure 2, HoPE first introduces a hybrid frequency allocation (HFA) strategy to better preserve the semantic preference property (Definition 4) in long-context modeling. Under this strategy, spatial information will be encoded with higher frequencies to capture local semantics, while the lowest frequencies will be set to zero (as in NoPE [25]) to facilitate long-range semantic modeling. Second, HoPE develops a dynamic temporal scaling (DTS) mechanism to enhance VLMs’ robustness to various video speeds and enable flexible inference under diverse context lengths. We will detail these strategies as follows:
@@ 4.1 Hybrid Frequency Allocation Strategy
To extend vanilla RoPE to multimodal scenarios, a common approach is to allocate different frequencies to encode different positional components $(t,x,y)$. For example, M-RoPE [2] assigns the highest frequencies for temporal modeling and lower frequencies for spatial encoding. In contrast, VideoRoPE [23] allocates the lowest frequencies for temporal modeling and achieves better empirical results. However, in Theorem 3.1, we theoretically prove that despite using lower frequencies is more ideal for semantic modeling, none of these frequency allocation strategies can maintain the ideal semantic preference property (Definition 4) over extended contexts.
To provide a stronger theoretical guarantee for the semantic preference property, we propose a hybrid frequency allocation strategy. As shown in Figure 1, we encode spatial information $(x,y)$ with high frequencies, as high frequencies are more sensitive to positional differences and thereby better at capturing local semantics [23, 32]. Following existing work [23], $x$ and $y$ are encoded in an interleaved manner to prevent biased spatial encoding. More importantly, unlike existing methods [18, 23, 2], we directly set the lowest frequencies to zero (as in NoPE [25]) to provide stronger guarantee for the semantic preference property (Definition 4), as shown in Figure 2. Specifically, for $d=128$, we interleave $x$ and $y$ positions in the first $96$ dimension of the rotation matrix and set the frequencies in the remaining $32$ dimensions to zero, which corresponds to an identity matrix:
$R_{\Delta x,\Delta y}=\operatorname{diag}(\begin{pmatrix}\cos\theta_{0}\Delta x&-\sin\theta_{0}\Delta x&0&0&\cdots&0&0&0&0\\ \sin\theta_{0}\Delta x&\cos\theta_{0}\Delta x&0&0&\cdots&0&0&0&0\\ 0&0&\cos\theta_{1}\Delta y&-\sin\theta_{1}\Delta y&\cdots&0&0&0&0\\ 0&0&\sin\theta_{1}\Delta y&\cos\theta_{1}\Delta y&\cdots&0&0&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&0&\dots&\cos\theta_{46}\Delta x&-\sin\theta_{46}\Delta x&0&0\\ 0&0&0&0&\dots&\sin\theta_{46}\Delta x&\cos\theta_{46}\Delta x&0&0\\ 0&0&0&0&\dots&0&0&\cos\theta_{47}\Delta y&-\sin\theta_{47}\Delta y\\ 0&0&0&0&\dots&0&0&\sin\theta_{47}\Delta y&\cos\theta_{47}\Delta y\\ \end{pmatrix},I_{32})$
We now provide a theoretical analysis of how this hybrid strategy helps the attention mechanism to capture long-range semantic similarities. Specifically, with Definition 4 and Theorem 3.1, we first formalize the condition under which semantic preference is preserved in multimodal RoPE.
Lemma 4.1 (Necessary Condition for Semantic Preference).
For a multimodal RoPE with rotation matrix $\mathbf{R}_{t,x,y}=\operatorname{diag}(\mathbf{R}_{t},\mathbf{R}_{x},\mathbf{R}_{y})$ (where $t,x,y$ do not denote their sequential order in the matrix), by Theorem 3.1, the semantic preference property (Definition 4) holds if the following condition is satisfied for all possible relative distances:
where $\sigma^{2}$ denotes the variance of each component in the query/key vector, $i_{t},i_{x},i_{y}$ denotes dimensions allocated to $t$, $x$, and $y$, respectively, and $\Delta t\in\{0,1,\dots,L-1\}$, $\Delta x\in\{0,1,\dots,H\}$, $\Delta y\in\{0,1,\dots,W\}$ represent the relative distances in these dimensions.
@@
Figure 2: Multimodal RoPEs use different frequencies for temporal modeling. M-RoPE uses the highest frequencies which are suboptimal for long-context modeling. VideoRoPE utilizes the lowest frequencies for more stable semantic modeling. Our HoPE, by using zero frequencies for temporal modeling, establishes the upper bound for semantic modeling capabilities across all strategies.
This lemma establishes a clear theoretical criterion for maintaining semantic preference with multimodal RoPE. It directly follows from our analysis in Theorem 3.1 and Appendix A.2, providing the foundation for our proposed method. Based on this condition, we now prove how our hybrid frequency allocation strategy provides stronger guarantees for semantic preference property over extended context. Specifically, HFA set $\theta_{i}=0$ for all $i\in i_{t}$. Hence, the temporal terms in Equation 5 reduce to $\sum_{i\in i_{t}}2\sigma^{2}\cdot 1$, noting that $\sum_{i\in i_{t}}2\sigma^{2}\cdot 1\geq\sum_{i\in i_{t}}2\sigma^{2}\cos(\Delta t\cdot\theta_{i})$ holds for any alternative choice of temporal frequencies $\theta_{i}$. Adding the identical spatial terms on both sides, we obtain:
@@
For multimodal position embeddings with dimensions allocated across temporal ($i_{t}$), and spatial dimensions ($i_{x},i_{y}$), setting $\theta_{i}=0$ for all temporal dimensions $i\in i_{t}$ maximizes the semantic preference guarantee in Definition 4 compared to any alternative frequency allocation strategy, particularly for extended context lengths.
Another interesting finding is that, if we set $|i_{t}|=d/4,|i_{x}|=|i_{y}|=d/8$ and $\theta_{i}=0,i\in i_{t}$, for any context length $t$ and spatial size $x,y$, semantic preference property invariably holds, as Equation 5 reduces to $\sum_{i=0}^{d/8-1}2\sigma^{2}(2+\mathrm{cos}(\Delta x\cdot\theta_{2i})+\mathrm{cos}(\Delta y\cdot\theta_{2i+1}))\geq 0.$ However, the empirical results of this approach are inferior to our proposed HoPE, probably due to the decreased number of frequencies for spatial modeling. More discussions are provided in Appendix B.4.

### v2

# HoPE: Hybrid of Position Embedding for Long Context Vision-Language Models

@@ Abstract
Vision-Language Models (VLMs) have made significant progress in multimodal tasks. However, their performance often deteriorates in long-context scenarios, particularly long videos. While Rotary Position Embedding (RoPE) has been widely adopted for length generalization in Large Language Models (LLMs), extending vanilla RoPE to capture the intricate spatial-temporal dependencies in videos remains an unsolved challenge. Existing methods typically allocate different frequencies within RoPE to encode 3D positional information. However, these allocation strategies mainly rely on heuristics, lacking in-depth theoretical analysis. In this paper, we first study how different allocation strategies impact the long-context capabilities of VLMs. Our analysis reveals that current multimodal RoPEs fail to reliably capture semantic similarities over extended contexts. To address this issue, we propose HoPE, a Hybrid of Position Embedding designed to improve the long-context capabilities of VLMs. HoPE introduces a hybrid frequency allocation strategy for reliable semantic modeling over arbitrarily long contexts, and a dynamic temporal scaling mechanism to facilitate robust learning and flexible inference across diverse context lengths. Extensive experiments across four video benchmarks on long video understanding and retrieval tasks demonstrate that HoPE consistently outperforms existing methods, confirming its effectiveness. Our code is available at https://github.com/hrlics/HoPE.
@@ 1 Introduction
Vision-Language Models (VLMs) [1, 2, 3, 4, 5] have achieved remarkable success in multimodal tasks, including visual question answering [6, 7, 8, 9], image captioning [10, 11], cross-modal retrieval [12, 13], and more [14, 15, 16]. However, VLMs suffer from significant performance degradation in long-context scenarios, particularly long videos [17, 18, 19, 20]. In such settings, VLMs even struggle with simple tasks like object counting and temporal localization [21, 22], revealing a critical limitation in their ability to model extended spatial-temporal dependencies. This limitation substantially hinders their real-world deployment, where input length often exceeds the context window they have been pretrained on.
Rotary Position Embedding (RoPE) [23] has been widely adopted for length generalization in text-based LLMs [24, 25, 26]. Specifically, RoPE incorporates positional information by partitioning the query and key vectors into 2-dimensional pairs and rotating each pair at a unique frequency that decreases as the dimensional index increases. Despite its advantages, directly applying 1D RoPE fails to capture the intricate spatial-temporal dependencies in videos. Several methods have been proposed to extend 1D RoPE for multimodal inputs [2, 27, 28]. Among these, the most common approach is to allocate different frequencies to encode different positional components, as shown in the upper plots of Figure 1. For example, M-RoPE [2] allocates the highest frequencies for temporal modeling ($t$), and the remaining low frequencies for spatial modeling ($x,y$). In contrast, VideoRoPE [28] proposes to allocate the lowest frequencies for temporal dimensions ($t$), and further applies a fixed scaling factor to scale the temporal indices of visual tokens, as shown in the lower plots of Figure 1. Despite their improved performance, two significant challenges remain unsolved. Firstly, current methods mainly rely on heuristics rather than theoretical analysis to determine the frequency allocation strategy. Second, applying a fixed and unidirectional scaling factor for all videos is suboptimal in real-world scenarios, where videos proceed at different speeds and demonstrate significant variance in information densities.
@@
To address these challenges, we begin with an in-depth theoretical analysis in Section 3, outlining the ideal properties that a multimodal RoPE should possess. Our analysis reveals that: (1) the flattening operation in vanilla RoPE inherently violates spatial-temporal locality prior, which is crucial in video modeling; (2) despite diverse frequency allocation strategies, existing multimodal RoPE variants fail to reliably capture semantic similarities over extended contexts; (3) temporal scaling of visual tokens should include both compression and expansion to accommodate varying video speeds in real-world scenarios. Guided by these insights, we propose HoPE, a Hybrid of Position Embedding designed to improve the long-context capabilities of VLMs. As shown in Figure 1, HoPE first introduces a hybrid frequency allocation strategy to facilitate long-range semantic modeling. In this strategy, higher frequencies, which are more sensitive to positional differences and better at capturing local features, are allocated to spatial components ($x,y$) in an interleaved manner. Meanwhile, the lower frequencies are directly set to zero and allocated to temporal component ($t$) to enable reliable semantic modeling. Moreover, HoPE develops a dynamic temporal scaling mechanism for lengthy visual tokens. This mechanism not only enhances VLMs’ robustness to various video speeds, which are common in real-world scenarios, but also offers flexible scaling during inference across diverse context lengths.
We summarize our contributions as follows:
  • To our best knowledge, we provide the first theoretical analysis of how different frequency allocation strategies in multimodal RoPEs impact the long-context capabilities of VLMs, offering insights for the design and analysis of future multimodal RoPEs.
  • Guided by our analysis, we propose HoPE, which consists of a hybrid frequency allocation strategy for reliable semantic modeling in long-context scenarios, and a dynamic temporal scaling mechanism for robust and flexible temporal comprehension.
  • Extensive experiments on four video benchmarks demonstrate that HoPE consistently outperforms existing RoPE variants, achieving improvements of 22.23% and 8.35% on long video retrieval and long video understanding tasks, confirming its effectiveness.
@@ 2 Preliminaries
$\small\mathbf{R}_{n-m}=\begin{pmatrix}\cos\theta_{0}(n-m)&-\sin\theta_{0}(n-m)&\cdots&0&0\\ \sin\theta_{0}(n-m)&\cos\theta_{0}(n-m)&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\cos\theta_{(d/2-1)}(n-m)&-\sin\theta_{(d/2-1)}(n-m)\\ 0&0&\cdots&\sin\theta_{(d/2-1)}(n-m)&\cos\theta_{(d/2-1)}(n-m)\end{pmatrix}$. (3)
It can be observed that through pairwise attention computation, the final rotation matrix naturally incorporates the relative position information (nm)(n-m) between the query-key pair.
@@ 3 Analysis
In this section, we conduct a comprehensive theoretical analysis of multimodal RoPE variants, aiming to answer the following questions: (1) Is vanilla RoPE enough for long-context VLMs? (2) How do different frequency allocation strategies impact semantic modeling in long-range multimodal contexts? (3) How should we assign the temporal indices for text and visual tokens?
@@ 3.1 Vanilla RoPE Fails in Spatial-Temporal Structure
Several recent VLMs [1, 5, 20, 33, 34, 35] still use vanilla RoPE for multimodal inputs. In their approach, each video frame is first encoded by a vision encoder (e.g., ViT [36]) and then flattened into a sequence of patch-level tokens. These visual tokens will be treated equally as text tokens for positional encoding, with each token incorporating only 1D temporal information. We show in Proposition 3.1 that this approach, while easy to implement, distorts spatial-temporal localities and fundamentally limits VLMs’ ability to model extended spatial-temporal dependencies.
@@
Conclusion 1. Directly applying vanilla RoPE to multimodal long-context inputs inherently fails to capture their complex spatial-temporal dependencies.
@@ 3.2 Current Multimodal RoPEs Are Unreliable in Long-Range Semantic Modeling
To capture the spatial-temporal structure of multimodal inputs, a recent VLM, Qwen2-VL [2], has introduced a Multimodal Rotary Position Embedding (M-RoPE). Concretely, M-RoPE partitions the $128$-dimensional rotary embedding into three distinct groups: the first $32$ dimensions for temporal information ($t$), the subsequent $48$ dimensions for horizontal spatial information ($x$), and the final $48$ dimensions for vertical spatial information ($y$), i.e., $\mathbf{R}_{t,x,y}=\operatorname{diag}(\mathbf{R}_{t},\mathbf{R}_{x},\mathbf{R}_{y})$. While this approach realizes a naive extension for RoPE, a fundamental question remains to be answered:
@@
This question arises from the fact that in RoPE, different dimensions carry unique frequencies ($\theta_{i}=b^{-2i/d}$, $i\in\{0,1,\dots,d/2-1\}$), as shown in Equation 3. Therefore, different strategies exist for frequency allocation in multimodal RoPE. As shown in Figure 1, M-RoPE allocates the highest frequencies for $t$, intermediate frequencies for $x$, and the lowest frequencies for $y$. In contrast, VideoRoPE [28] proposes to assign the lowest frequencies to temporal modeling $(t)$ and high frequencies to spatial dimensions $(x,y)$. Their empirical justification stems from attention pattern analysis, which reveals that dimensions encoded with the lowest frequencies exhibit a more pronounced attention sink phenomenon [37], which has proven to be effective in long-context modeling. However, we argue that using the lowest frequencies for temporal modeling is still unreliable in capturing semantic similarities in extended multimodal contexts. Specifically, we first introduce semantic preference, a property where attention mechanisms should prioritize semantically similar tokens regardless of their relative distance, and formally define this concept in Definition 3.1.
@@
$\mathbb{E}_{\mathbf{q},\mathbf{k},\delta}[\mathbf{q}\mathbf{R}_{\Delta t\Delta x\Delta y}\mathbf{k}^{\prime\top}-\mathbf{q}\mathbf{R}_{\Delta t\Delta x\Delta y}\mathbf{k}^{\top}]\geq 0,$
@@
Let $X=[x_{1},x_{2},\dots,x_{L}]$ be an input sequence, and let RoPE use any fixed set of temporal frequencies (e.g., highest or lowest). Then there exists a critical length $L_{c}$ such that for all $L\geq L_{c}$, the semantic preference property (Definition 3.1) is violated.
Conclusion 2. There exist different frequency allocation strategies to extend vanilla RoPE to multimodal RoPE. However, we prove that none of these strategies can reliably maintain the semantic preference property over a sufficiently long context.
3.3 How to Assign Positional Index for Multimodal Inputs?
Currently, most VLMs [1, 2, 3, 4, 19, 20] adopt the same temporal stride for video frames and text tokens, as shown in Figure 1. However, this approach overlooks the inherent difference in information densities between text and visual tokens. To address this issue, VideoRoPE [28] applies a fixed scaling factor (implemented as $2$) to adjust the temporal indices of visual tokens, achieving better empirical performance. However, this rigid scaling approach lacks the flexibility needed for diverse real-world videos, which naturally vary in pace and information density. A more ideal approach would incorporate both temporal compression and expansion capabilities, allowing the model to learn multi-scale temporal relationships, thereby enabling more robust temporal modeling.
Conclusion 3. Temporal index scaling of visual tokens is crucial for balancing multimodal information, yet current methods lack flexibility and bidirectionality.
4 HoPE: Hybrid of Position Embedding for Long Context VLMs
To address the above challenges, we propose HoPE, a Hybrid of Position Embedding designed to improve the long-context capability of VLMs. As illustrated in Figure 1 and Figure 2, HoPE first introduces a hybrid frequency allocation (HFA) strategy to better preserve the semantic preference property (Definition 3.1) in long-context modeling. Under this strategy, spatial information will be encoded with higher frequencies to capture local semantics, while the lowest frequencies will be set to zero (as in NoPE [30]) to facilitate long-range semantic modeling. Second, HoPE develops a dynamic temporal scaling (DTS) mechanism to enhance VLMs’ robustness to various video speeds and enable flexible inference under diverse context lengths. We will detail these strategies as follows:
@@ 4.1 Hybrid Frequency Allocation Strategy
To extend vanilla RoPE to multimodal scenarios, a common approach is to allocate different frequencies to encode different positional components $(t,x,y)$. For example, M-RoPE [2] assigns the highest frequencies for temporal modeling and lower frequencies for spatial encoding. In contrast, VideoRoPE [28] allocates the lowest frequencies for temporal modeling, achieving better empirical results. However, in Theorem 3.1, we theoretically prove that, despite using lower frequencies being more ideal for semantic modeling, none of these frequency allocation strategies can maintain the ideal semantic preference property (Definition 3.1) over extended contexts.
To provide a stronger theoretical guarantee for the semantic preference property, we propose a hybrid frequency allocation strategy. As shown in Figure 1, we encode spatial information $(x,y)$ with high frequencies, as high frequencies are more sensitive to positional differences and thereby better at capturing local semantics [28, 38]. Following existing work [28], $x$ and $y$ are encoded in an interleaved manner to prevent biased spatial encoding. More importantly, unlike existing methods [23, 28, 2], we directly set the lowest frequencies to zero (as in NoPE [30]) to provide a stronger guarantee for the semantic preference property (Definition 3.1), as shown in Figure 2. Specifically, for $d=128$, we interleave $x$ and $y$ positions in the first $96$ dimensions of the rotation matrix and set the frequencies in the remaining $32$ dimensions to zero, which corresponds to an identity matrix:
$\tiny R_{\Delta x,\Delta y}=\operatorname{diag}(\begin{pmatrix}\cos\theta_{0}\Delta x&-\sin\theta_{0}\Delta x&0&0&\cdots&0&0&0&0\\ \sin\theta_{0}\Delta x&\cos\theta_{0}\Delta x&0&0&\cdots&0&0&0&0\\ 0&0&\cos\theta_{1}\Delta y&-\sin\theta_{1}\Delta y&\cdots&0&0&0&0\\ 0&0&\sin\theta_{1}\Delta y&\cos\theta_{1}\Delta y&\cdots&0&0&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&0&\dots&\cos\theta_{46}\Delta x&-\sin\theta_{46}\Delta x&0&0\\ 0&0&0&0&\dots&\sin\theta_{46}\Delta x&\cos\theta_{46}\Delta x&0&0\\ 0&0&0&0&\dots&0&0&\cos\theta_{47}\Delta y&-\sin\theta_{47}\Delta y\\ 0&0&0&0&\dots&0&0&\sin\theta_{47}\Delta y&\cos\theta_{47}\Delta y\\ \end{pmatrix},I_{32})$
We now provide a theoretical analysis of how this hybrid strategy helps the attention mechanism to capture long-range semantic similarities. Building on Definition 3.1 and Theorem 3.1, we first formalize the condition under which semantic preference is preserved in multimodal RoPE.
@@
Figure 2: Multimodal RoPEs use different frequencies for temporal modeling. M-RoPE uses the highest frequencies, which are suboptimal for long-context modeling. VideoRoPE utilizes the lowest frequencies for more stable semantic modeling. Our HoPE, employing zero frequencies for temporal modeling, establishes the upper bound of semantic modeling capabilities across all strategies.
In particular, Lemma 4.1 establishes a clear theoretical criterion for maintaining semantic preference with multimodal RoPE. It directly follows from our analysis in Theorem 3.1 and Appendix A.2, providing the theoretical foundation for our proposed method.
Lemma 4.1 (Necessary Condition for Semantic Preference).
For a multimodal RoPE with rotation matrix $\mathbf{R}_{t,x,y}=\operatorname{diag}(\mathbf{R}_{t},\mathbf{R}_{x},\mathbf{R}_{y})$, the semantic preference property (Definition 3.1) holds if, for all possible relative distances, $\sum_{i\in i_{t}}2\sigma^{2}\mathrm{cos}(\Delta t\cdot\theta_{i})+\sum_{i\in i_{x}}2\sigma^{2}\mathrm{cos}(\Delta x\cdot\theta_{i})+\sum_{i\in i_{y}}2\sigma^{2}\mathrm{cos}(\Delta y\cdot\theta_{i})\geq 0,$ where $\sigma^{2}$ is the variance of each component in the query/key vector, $i_{t},i_{x},i_{y}$ are dimensions allocated to $t,x,y$, and $\Delta t\in\{0,1,\dots,L-1\}$, $\Delta x\in\{0,1,\dots,H\}$, $\Delta y\in\{0,1,\dots,W\}$.
Based on this Lemma, we now prove how our hybrid frequency allocation strategy provides stronger guarantees for the semantic preference property over extended contexts. Specifically, HFA set $\theta_{i}=0$ for all $i\in i_{t}$. Hence, the temporal terms in Lemma 4.1 reduce to $\sum_{i\in i_{t}}2\sigma^{2}\cdot 1$, noting that $\sum_{i\in i_{t}}2\sigma^{2}\cdot 1\geq\sum_{i\in i_{t}}2\sigma^{2}\cos(\Delta t\cdot\theta_{i})$ holds for any choice of temporal frequencies $\theta_{i}$. Adding the identical spatial terms on both sides, we obtain:
@@
For multimodal position embeddings with dimensions allocated across temporal ($i_{t}$), and spatial components ($i_{x},i_{y}$), setting $\theta_{i}=0$ for all temporal dimensions $i\in i_{t}$ maximizes the semantic preference guarantee in Definition 3.1, compared to any alternative frequency allocation strategy, particularly under extended context lengths.
Another interesting finding is that, if we set $|i_{t}|=d/4,|i_{x}|=|i_{y}|=d/8$ and $\theta_{i}=0,i\in i_{t}$, for any context length $t$ and spatial size $x,y$, semantic preference property invariably holds, as Lemma 4.1 reduces to $\sum_{i=0}^{d/8-1}2\sigma^{2}(2+\mathrm{cos}(\Delta x\cdot\theta_{2i})+\mathrm{cos}(\Delta y\cdot\theta_{2i+1}))\geq 0.$ However, the empirical results of this approach are inferior to our proposed HoPE, probably due to the decreased number of frequencies for spatial modeling. More discussions are provided in Appendix B.3.